# Automated Detection of Visual Attribute Reliance with a Self-Reflective Agent

### v1

@@ 5.3 Revealing attribute reliance in pretrained vision models
We deploy SAIA to detect attribute reliances in two pre-trained vision models: the CLIP-ViT image encoder (Radford et al., 2021) trained to align image and text representations, and the YOLOv8 model (Jocher et al., 2023) trained for object detection in autonomous driving settings. With CLIP, we perform object recognition by measuring the cosine similarity of the image with a target prompt (e.g., “A picture of a scientist”). For YOLOv8 we measure the detection score of the target object class. Figures 1 and 7 show that SAIA can generate natural-language descriptions of attribute reliance in various contexts. The generated descriptions are shown to be predictive of model behavior, as model scores increase when the reliance is satisfied and decrease when it is absent. Surprisingly, SAIA reveals dependencies that were never observed before, such as the reliance of clip on traditional laboratory settings when detecting scientist, and YOLOv8 dependency on bikers’ poses. We note that SAIA’s goal is to surface such dependencies rather than to assess their desirability or harm, allowing practitioners to make informed judgments based on specific downstream use cases.
@@ Acknowledgments
We are grateful for the support of ARL grant #W911NF-24-2-0069, MIT-IBM Watson AI Lab grant #W1771646, Hyundai NGV, ONR MURI grant #033697-00007, and PID2022-138721NB-I00, funded by MCIN/AEI/10.13039/501100011033 and FEDER, EU.
@@ A.4 SAIA API and implementation details
SAIA’s API is based on that of MAIA Rott Shaham et al. [2024] with a few important modifications: (i) SAIA is able to define experiments in free-form code blocks, where initially MAIA had to define all its code within a function (execute_command) that was then executed by the codebase. This allows SAIA to both write multiple blocks of code per experiment and to access variables defined in previous experiments. (ii) SAIA is now able to log and display any text/images generated during its experiments in the format of its choosing with a single, flexible function (display), whereas MAIA had to rely on individual tools to display their own results in predetermined formats. (iii) SAIA can use more recent VLLMs, particularly Claude-3.5-sonnet, where the original codebase used gpt-4-vision-preview.
@@ A.5 Agent tools
    • Flux Image Generator Labs [2024]
    • DALL·E 3 OpenAI [2023] (OpenAI API) (used for the evaluation)
@@
    • InstructDiffusion (Stable Diffusion variant with instruction tuning) Geng et al. [2023]
@@
    • GPT-4o OpenAI [2024], used via API to describe image regions and summarize visual commonalities across images.

### v2

@@ 5.3 Revealing attribute reliance in pretrained vision models
We deploy SAIA to detect attribute reliances in two pre-trained vision models: the CLIP-ViT image encoder (radford2021learning) trained to align image and text representations, and the YOLOv8 model (yolov8_ultralytics) trained for object detection in autonomous driving settings. With CLIP, we perform object recognition by measuring the cosine similarity of the image with a target prompt (e.g., “A picture of a scientist”). For YOLOv8 we measure the detection score of the target object class. Figures 1 and 7 show that SAIA can generate natural-language descriptions of attribute reliance in various contexts. The generated descriptions are shown to be predictive of model behavior, as model scores increase when the reliance is satisfied and decrease when it is absent. Surprisingly, SAIA reveals dependencies that were never observed before, such as the reliance of clip on traditional laboratory settings when detecting scientist, and YOLOv8 dependency on bikers’ poses. We note that SAIA’s goal is to surface such dependencies rather than to assess their desirability or harm, allowing practitioners to make informed judgments based on specific downstream use cases.
@@ Acknowledgments
We are grateful for the support of ARL grant #W911NF-24-2-0069, MIT-IBM Watson AI Lab grant #W1771646, Hyundai Motor Company, ONR MURI grant #033697-00007, and PID2022-138721NB-I00, funded by MCIN/AEI/10.13039/501100011033 and FEDER, EU.
@@ A.4 SAIA API and implementation details
SAIA’s API is based on that of MAIA shaham2024multimodal with a few important modifications: (i) SAIA is able to define experiments in free-form code blocks, where initially MAIA had to define all its code within a function (execute_command) that was then executed by the codebase. This allows SAIA to both write multiple blocks of code per experiment and to access variables defined in previous experiments. (ii) SAIA is now able to log and display any text/images generated during its experiments in the format of its choosing with a single, flexible function (display), whereas MAIA had to rely on individual tools to display their own results in predetermined formats. (iii) SAIA can use more recent VLLMs, particularly Claude-3.5-sonnet, where the original codebase used gpt-4-vision-preview.
@@ A.5 Agent tools
    • Flux Image Generator flux2024
    • DALL·E 3 openai2023dalle3 (OpenAI API) (used for the evaluation)
@@
    • InstructDiffusion (Stable Diffusion variant with instruction tuning) Geng23instructdiff
@@
    • GPT-4o openai2024gpt4o, used via API to describe image regions and summarize visual commonalities across images.